# Arcee’s MergeKit: A Toolkit for Merging Large Language Models

### 20 Mar 2024 v1

@@ Abstract
The rapid expansion of the open-source language model landscape presents an opportunity to merge the competencies of these model checkpoints by combining their parameters. Advances in transfer learning, the process of fine-tuning pre-trained models for specific tasks, has resulted in the development of vast amounts of task-specific models, typically specialized in individual tasks and unable to utilize each other’s strengths. Model merging facilitates the creation of multitask models without the need for additional training, offering a promising avenue for enhancing model performance and versatility. By preserving the intrinsic capabilities of the original models, model merging addresses complex challenges in AI - including the difficulties of catastrophic forgetting and multi-task learning. To support this expanding area of research, we introduce MergeKit, a comprehensive, open-source library designed to facilitate the application of model merging strategies. MergeKit offers an extensible framework to efficiently merge models on any hardware, providing utility to researchers and practitioners. To date, thousands of models have been merged by the open-source community, leading to the creation of some of the worlds most powerful open-source model checkpoints, as assessed by the Open LLM Leaderboard. The library is accessible at https://github.com/arcee-ai/MergeKit.
@@ 1 Introduction
Over the last year, we noticed a rapid development of open-source LLM models and these LLMs are accessible via the Hugging Face model hub Wolf et al. (2019). These models are typically trained on a corpus comprising trillions of tokens and they consist of parameters in the range of 1-70 billions Minaee et al. (2024); Zhang et al. (2024). The landscape of open-source checkpoints is rich and varied, with a distinction generally made between pre-trained checkpoints Zhuang et al. (2020) and those specifically aligned for instruction-following tasks across diverse domains (e.g., coding Roziere et al. (2023), medical applications Wu et al. (2023), etc.). However, fine-tuning a separate model for each task raises two major challenges: (1) For each new task, the task-specific model should be stored and deployed separately, and (2) models trained independently cannot utilize insights from related tasks to enhance performance within their domain or generalize beyond it Sanh et al. (2021); Ramé et al. (2023); Yadav et al. (2024); Yu et al. (2023).
Training these models from scratch represents a formidable investment, exemplified by the Mistral-7B model Jiang et al. (2023), which demands an outlay of 2 to 3 million USD. In addition, further fine-tuning pre-trained models can lead to catastrophic forgetting De Lange et al. (2021), which is a phenomenon occurring when further refinement of these models results in a degradation of their original general capabilities, thereby impeding their ability to perform optimally across a range of tasks Cheng et al. (2023); Wu et al. (2024). Furthermore, the process of aligning these models to respond in a favorable fashion takes extensive effort to gather human preference data, and is impossible for most teams to replicate Wang et al. (2023); Rafailov et al. (2024). This backdrop sets the stage for the pivotal question of how to maximally leverage existing pretrained checkpoints for research and industrial use cases. In this context, research in model merging has emerged as a transformative strategy, which combines the parameters of multiple individual models, often trained for specific tasks, into a single unified model. It enables multi-task learning and continual learning while reducing the risk of catastrophic forgetting, all without the prohibitive costs of re-training from scratch Yadav et al. (2023); Yu et al. (2023).
@@ 2.1 The Concept of Model Merging
Model merging Ainsworth et al. (2022), though a relatively recent focal point within the research community, builds upon a foundation laid by numerous prior studies. At its core, model merging involves integrating two or more pre-trained models — whether they’ve been trained on identical tasks or distinct architectures — into a unified model that retains the strengths and capabilities of all the original models. Much of this builds upon concepts explored with weight averaging, such as in Utans (1996). The success of these techniques relies on the concept of mode connectivity Garipov et al. (2018). In the simplest cases, techniques take advantage of the Linear Mode Connectivity (LMC) Entezari et al. (2021) of models fine-tuned from a common pre-trained model Nagarajan and Kolter (2019); Neyshabur et al. (2021). Other works build on this concept by employing the permutation symmetry and applying transformations to the weights of models that bring them into common regions in the loss landscape Ainsworth et al. (2022); Stoica et al. (2023); Verma and Elbayad (2024).
@@ 2.2.1 Merging Models with Both Identical Architectures and Initializations
It should be emphasized that the above methods described in this section require no training data for merging operations nor further fine-tuning post-merging. Compared to the methods mentioned above, other approaches expand on this paradigm by introducing varied weightings for different neural network parameters, requiring the computation of activations with training data to determine these varying weightings. Matena and Raffel (2022) explore the use of the Fisher information matrix and Jin et al. (2022) introduce the RegMean Jin et al. (2022) method, which allows merges to produce optimal weights with respect to $L2$ distance to model predictions while keeping training data private.
@@ 2.2.2 Merging Models with Identical Architectures and Different Initializations
This section explores merging methods that extend beyond simply combining checkpoints with identical initializations. Previous research has indicated that when dealing with checkpoints from different initializations, the straightforward Linear Model Combination (LMC) approach falls short Ainsworth et al. (2022). So basically this line of work utilizes the permutation symmetry of the different initializations of checkpoints prior to the merging Verma and Elbayad (2024); Ainsworth et al. (2022)
@@
It is important to note that the above methodologies facilitate merging of models which share architectures and sizes, despite them not being linearly mode connected — differences that may stem from varied random initializations, extensive continued pre-training, or other distinct reasons.
Further expanding the scope, compared to above-mentioned methods, ZipIt Stoica et al. (2023) explores the possibility of merging models of similar architectures that have been trained on distinct tasks. This method generalizes model merging by supporting the ‘zipping’ of correlated features within and across each model, and it permits partial merging of models up to a specified layer, thereby producing a multi-head model. This represents a significant step forward in the flexibility and applicability of model merging techniques, addressing the challenge of preserving and integrating the knowledge of models from different domains within a single unified framework without any additional training.
@@ 2.2.3 Fusing Models with Different Architectures
While not strictly model merging, Composition to Augment Language Models (CALM) Bansal et al. (2024) and recent approaches like knowledge fusion approach for large language models (FUSELLM) Wan et al. (2024) represent significant steps towards the fusion of models with diverse architectures. CALM utilizes cross-attention mechanisms to blend representations from different models, aiming to leverage their combined strengths and functionalities, thus facilitating integration across varied neural network structures. Similarly, FUSELLM Wan et al. (2024) focuses on harnessing the generative capabilities of source Large Language Models (LLMs) by aligning and fusing their probabilistic distributions. This strategy aims to externalize and amplify the collective knowledge and unique advantages of individual models, enhancing the overall capabilities of the fused LLM. In contrast to the approaches discussed above, the methods outlined in this section require additional continual pre-training of the models.
Figure 1: Classification of model merging methods. We currently support the model merging methods outlined on the left, and we are actively working to incorporate additional merging techniques such as ZipIt, OT Fusion, and Git Rebasing.
@@ 2.3 Practical Use Cases of Model Merging
Model merging has found its place in a variety of practical applications, significantly impacting the landscape of machine learning models available on platforms such as HuggingFace’s model hub Wolf et al. (2019). These merged models, which will be detailed further, have demonstrated competitive performance across a range of tasks. A notable example of this is BioMistral Labrak et al. (2024), a project that merges domain-adapted checkpoints with existing Mistral chat variants, showcasing the efficacy of model merging in enhancing performance in specialized domains. OpenPipe’s Mistral 7B Fine-Tune Optimized Corbitt (2023) demonstrates the promise of merging fine-tuned models to produce a high-quality base for further tuning, as in Choshen et al. (2022). Wei et al. (2024) illustrate that employing the MergeKit tool for model fusion is a successful method for enhancing the performance of hallucination detection.
The success stories of merged models underscore the technique’s value in continuous learning and multi-task learning scenarios. By utilizing the broad spectrum of open-sourced LLMs, model merging enables the creation of versatile and robust models capable of excelling at multiple tasks simultaneously or adapting to new domains without the need for training from scratch. This approach not only maximizes the utility of existing resources but also paves the way for innovative solutions in leveraging pre-trained models for complex, real-world problems.
@@ 3 Library Design: Key Design Principles
Figure 2: MergeKit structure with key modules for adding new merge methods. The diagram depicts the workflow for introducing new merge methods in the MergeKit repository. Initially, tensors are extracted from two models, A and B, and processed by the ’Architecture’ module to ensure their structural compatibility. Next, the ’Plan’ component formulates a strategy detailing the merge process. This plan is then relayed to a ’Graph’, outlining the necessary operations for merging. During ’Graph Execution’, these operations are performed, resulting in the ’Merged Model’—the integration of Models A and B via a specified merging technique within the system’s framework.
@@ 3.2 Modularity: Plug-and-Play Components
MergeKit is designed with composability and reusability as guiding principles. Merge methods are designed to be interchangeable and easy to add. Components are structured such that they can be added, removed, or interchanged to allow customization and experimentation. Wherever possible, components are designed to be useful standalone for external use. For instance, MergeKit’s lazy tensor loading functionality is a core component of the toolkit but is also simple and convenient to pull into one-off scripts.
@@ 3.4 Scalability: Efficiency and Performance Optimization
MergeKit is designed specifically to address the challenge of merging large pre-trained language models, ensuring compatibility and efficiency across a diverse range of computational resources.. At the heart of its efficiency is an out-of-core approach to model merging. By loading only the tensors necessary for each individual operation into working memory, MergeKit can scale from a high-end research cluster all the way down to a personal laptop with no GPU and little RAM.
@@ 5 Popularity and Effectiveness of MergeKit
Figure 4: Performance comparison of top 50 models on Open LLM Leaderboard based on ‘Average’ score. Merged models are highlighted in pink. What does the dotted line say
@@
Table 1: The table compares the Llama2-7B Chat and Meditron-7B Chen et al. (2023) models, plus their merged variants, using MergeKit techniques across medical and general benchmarks. It highlights base and merged model performances, with the best-performing models in bold for each metric.
The utilization of model merging techniques in the development and refinement of LLMs has gained considerable attention within the machine learning community. Figure 3 shows the number of GitHub stars MergeKit has received over the last six months. This trend is evidenced by the Open LLM Leaderboard Beeching et al. (2023) data as of March 15th, 2024, which highlights the increasing prevalence of merged models among high-performing LLMs. Specifically, merged models represent 20% of the top 50 and 34% of the top 100 models, underscoring their significance in current LLM advancements. Plus, the current best-performing 3B555liminerity/Phigments12, 7B666liminerity/M7-7b, and the third top 13B [^7] RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2 open-source models on the Open LLM Leaderboard are merged models using MergeKit. In addition, the MergeKit repository’s meteoric rise in popularity is clearly illustrated by Figure 3, highlighting a significant and accelerating number of GitHub stars over time — an indicator of its growing influence and user endorsement in the developer community.
Figure 4 indicates a visual comparison of the performance of the top 50 open-source models ranked by their ‘Average’ scores on the Open LLM Leaderboard Park (2023). The ‘Average’ score is the mean value of evaluation scores from ‘ARC’Clark et al. (2018), ‘HellaSwag’Zellers et al. (2019), ‘MMLU’Hendrycks et al. (2021), ‘TruthfulQA’Lin et al. (2022), ‘Winogrande’Sakaguchi et al. (2019), ‘GSM8K’Cobbe et al. (2021) benchmarks. The line plot in black illustrates the parameter count for each model, and the bar chart represents the average score of each model. Merged models are depicted in pink bars, while the rest are colored in blue. The majority of leading merged models typically contain around 7 to 13 billion parameters. Among them, the top-performing merged model, ‘RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2’ [^8] https://huggingface.co/RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2, ranks 17th on the list. Remarkably, apart from models utilizing mixture of experts (MoE) Gale et al. (2023), majority of unmerged models surpassing this performance comprise approximately 35 to 70+ billion parameters. As depicted by the Open LLM Leaderboard, Merged models, particularly those incorporating mixture of experts (MoE) architecture, emerge as formidable contenders, offering superior performance while optimizing computational resources — a testament to the transformative potential of model merging in advancing the frontier of natural language processing.
@@ 6 Conclusion and Future Work
As stewards of the open-source community dedicated to the advancement of LLMs, our work with MergeKit underscores a commitment to democratizing access to cutting-edge AI technologies while fostering an environment of ethical integrity and continuous improvement. By providing an open-source toolkit that enables the merging of model checkpoints, we aim to enhance the collaborative capabilities of researchers, developers, and practitioners across the globe, encouraging innovation and the sharing of knowledge. In doing so, we are acutely aware of the necessity to uphold principles of fairness, accountability, and transparency within this community. This includes the proactive identification and mitigation of biases within merged models, ensuring the ethical use of data, and maintaining the privacy and security of information. Our commitment extends beyond technological advancements, encompassing the responsibility to engage with diverse stakeholders, gather feedback, and adapt our approaches to address ethical concerns effectively. We recognize the imperative to continually evolve our practices, striving for solutions that not only advance the field of AI but do so with an unwavering commitment to the betterment of society.

### 21 Mar 2024 v2

@@ Abstract
The rapid expansion of the open-source language model landscape presents an opportunity to merge the competencies of these model checkpoints by combining their parameters. Advances in transfer learning, the process of fine-tuning pretrained models for specific tasks, has resulted in the development of vast amounts of task-specific models, typically specialized in individual tasks and unable to utilize each other’s strengths. Model merging facilitates the creation of multitask models without the need for additional training, offering a promising avenue for enhancing model performance and versatility. By preserving the intrinsic capabilities of the original models, model merging addresses complex challenges in AI - including the difficulties of catastrophic forgetting and multitask learning. To support this expanding area of research, we introduce MergeKit, a comprehensive, open-source library designed to facilitate the application of model merging strategies. MergeKit offers an extensible framework to efficiently merge models on any hardware, providing utility to researchers and practitioners. To date, thousands of models have been merged by the open-source community, leading to the creation of some of the worlds most powerful open-source model checkpoints, as assessed by the Open LLM Leaderboard. The library is accessible at https://github.com/arcee-ai/MergeKit.
@@ 1 Introduction
Over the last year, we noticed a rapid development in open-source LLM models and these LLMs are accessible via the Hugging Face model hub Wolf et al. (2019). These models are typically trained on a corpus comprising trillions of tokens and they consist of parameters in the range of 1-70 billions Minaee et al. (2024); Zhang et al. (2024). The landscape of open-source checkpoints is rich and varied, with a distinction generally made between pretrained checkpoints Zhuang et al. (2020) and those specifically aligned for instruction-following tasks across diverse domains (e.g., coding Roziere et al. (2023), medical applications Wu et al. (2023), etc.). However, fine-tuning a separate model for each task raises two major challenges: (1) For each new task, the task-specific model should be stored and deployed separately, and (2) models trained independently cannot utilize insights from related tasks to enhance performance within their domain or generalize beyond it Sanh et al. (2021); Ramé et al. (2023); Yadav et al. (2024); Yu et al. (2023).
Training these models from scratch represents a formidable investment, exemplified by the Mistral-7B model Jiang et al. (2023), which demands an outlay of 2 to 3 million USD. In addition, further fine-tuning pretrained models can lead to catastrophic forgetting De Lange et al. (2021), which is a phenomenon occurring when further refinement of these models results in a degradation of their original general capabilities, thereby impeding their ability to perform optimally across a range of tasks Cheng et al. (2023); Wu et al. (2024). Furthermore, the process of aligning these models to respond in a favorable fashion takes extensive effort to gather human preference data, and is impossible for most teams to replicate Wang et al. (2023); Rafailov et al. (2024). This backdrop sets the stage for the pivotal question of how to maximally leverage existing pretrained checkpoints for research and industrial use cases. In this context, research in model merging has emerged as a transformative strategy, which combines the parameters of multiple individual models, often trained for specific tasks, into a single unified model. It enables multitask learning and continual learning while reducing the risk of catastrophic forgetting, all without the prohibitive costs of retraining from scratch Yadav et al. (2023); Yu et al. (2023).
@@ 2.1 The Concept of Model Merging
Model merging Ainsworth et al. (2022), though a relatively recent focal point within the research community, builds upon a foundation laid by numerous prior studies. At its core, model merging involves integrating two or more pretrained models — whether they’ve been trained on identical tasks or distinct architectures — into a unified model that retains the strengths and capabilities of all the original models. Much of this builds upon concepts explored with weight averaging, such as in Utans (1996). The success of these techniques relies on the concept of mode connectivity Garipov et al. (2018). In the simplest cases, techniques take advantage of the Linear Mode Connectivity (LMC) Entezari et al. (2021) of models fine-tuned from a common pretrained model Nagarajan and Kolter (2019); Neyshabur et al. (2021). Other works build on this concept by employing the permutation symmetry and applying transformations to the weights of models that bring them into common regions in the loss landscape Ainsworth et al. (2022); Stoica et al. (2023); Verma and Elbayad (2024).
@@ 2.2.1 Merging Models with Both Identical Architectures and Initializations
It should be emphasized that the methods described above do not require training data for merging operations or further fine-tuning post-merging. Compared to the methods mentioned above, other approaches expand on this paradigm by introducing varied weightings for different neural network parameters, requiring the computation of activations with training data to determine these varying weightings. Matena and Raffel (2022) explore the use of the Fisher information matrix and Jin et al. (2022) introduce the RegMean Jin et al. (2022) method, which allows merges to produce optimal weights with respect to $L2$ distance to model predictions while keeping training data private.
@@ 2.2.2 Merging Models with Identical Architectures and Different Initializations
This section explores merging methods that extend beyond simply combining checkpoints with identical initializations. Previous research has indicated that when dealing with checkpoints from different initializations, the straightforward Linear Model Combination (LMC) approach falls short Ainsworth et al. (2022). So, basically, this line of work utilizes the permutation symmetry of the different initializations of checkpoints prior to merging Verma and Elbayad (2024); Ainsworth et al. (2022)
@@
It is important to note that the above methodologies facilitate merging of models which share architectures and sizes, despite them not being linearly mode connected — differences that may stem from varied random initializations, extensive continued pretraining, or other distinct reasons.
Further expanding the scope, compared to above-mentioned methods, ZipIt Stoica et al. (2023) explores the possibility of merging models of similar architectures that have been trained on distinct tasks. This method generalizes model merging by supporting ‘zipping’ of correlated features within and across each model, and it permits partial merging of models up to a specified layer, thereby producing a multi-head model. This represents a significant step forward in the flexibility and applicability of model merging techniques, addressing the challenge of preserving and integrating the knowledge of models from different domains within a single unified framework without any additional training.
@@ 2.2.3 Fusing Models with Different Architectures
While not strictly model merging, Composition to Augment Language Models (CALM) Bansal et al. (2024) and recent approaches like knowledge fusion approach for large language models (FUSELLM) Wan et al. (2024) represent significant steps towards the fusion of models with diverse architectures. CALM utilizes cross-attention mechanisms to blend representations from different models, aiming to leverage their combined strengths and functionalities, thus facilitating integration across varied neural network structures. Similarly, FUSELLM Wan et al. (2024) focuses on harnessing the generative capabilities of source Large Language Models (LLMs) by aligning and fusing their probabilistic distributions. This strategy aims to externalize and amplify the collective knowledge and unique advantages of individual models, enhancing the overall capabilities of the fused LLM. In contrast to the approaches discussed above, the methods outlined in this section require additional continual pretraining of the models.
Figure 1: Classification of model merging methods. We currently support the model merging methods outlined on the left, and we are actively working to incorporate additional merging techniques such as ZipIt, OT Fusion, and Git Rebasin.
@@ 2.3 Practical Use Cases of Model Merging
Model merging has found its place in a variety of practical applications, significantly impacting the landscape of machine learning models available on platforms such as HuggingFace’s model hub Wolf et al. (2019). These merged models, which will be detailed further, have demonstrated competitive performance across a range of tasks. A notable example of this is BioMistral Labrak et al. (2024), a project that merges domain-adapted checkpoints with existing Mistral chat variants, showcasing the efficacy of model merging in enhancing performance in specialized domains. OpenPipe’s Mistral 7B Fine-Tune Optimized Corbitt (2023) demonstrates the promise of merging fine-tuned models to produce a high-quality base for further tuning as in Choshen et al. (2022). Wei et al. (2024) illustrate that employing the MergeKit tool for model fusion is a successful method for enhancing the performance of hallucination detection.
The success stories of merged models underscore the technique’s value in continuous learning and multitask learning scenarios. By utilizing the broad spectrum of open-sourced LLMs, model merging enables the creation of versatile and robust models capable of excelling at multiple tasks simultaneously or adapting to new domains without the need for training from scratch. This approach not only maximizes the utility of existing resources but also paves the way for innovative solutions in leveraging pretrained models for complex, real-world problems.
@@ 3 Library Design: Key Design Principles
Figure 2: MergeKit structure with key modules for adding new merge methods. The diagram depicts the workflow for introducing new merge methods in the MergeKit repository. Initially, tensors are extracted from two models, A and B, and processed by the ‘Architecture’ module to ensure their structural compatibility. Next, the ‘Plan’ component formulates a strategy detailing the merge process. This plan is then relayed to a ‘Graph’, outlining the necessary operations for merging. During ‘Graph Execution’, these operations are performed, resulting in the ‘Merged Model’—the integration of Models A and B via a specified merging technique within the system’s framework.
@@ 3.2 Modularity: Plug-and-Play Components
MergeKit is designed with composability and reusability as guiding principles. Merge methods are designed to be interchangeable and easy-to-add. Components are structured such that they can be added, removed, or interchanged to allow customization and experimentation. Wherever possible, components are designed to be useful standalone for external use. For instance, MergeKit’s lazy tensor loading functionality is a core component of the toolkit, but is also simple and convenient to pull into one-off scripts.
@@ 3.4 Scalability: Efficiency and Performance Optimization
MergeKit is designed specifically to address the challenge of merging large pretrained language models, ensuring compatibility and efficiency across a diverse range of computational resources.. At the heart of its efficiency is an out-of-core approach to model merging. By loading only the tensors necessary for each individual operation into working memory, MergeKit can scale from a high-end research cluster all the way down to a personal laptop with no GPU and limited RAM.
@@ 5 Popularity and Effectiveness of MergeKit
Figure 4: Performance comparison of top 50 models on Open LLM Leaderboard based on ‘Average’ score. Merged models are highlighted in ruby red color. The dotted line represents the number of parameters in each model.
@@
Table 1: Comparison of the Llama2-7B Chat and Meditron-7B Chen et al. (2023) models, plus their merged variants, using MergeKit techniques across medical and general benchmarks. It highlights the best-performing models in bold for each metric.
The utilization of model merging techniques in the development and refinement of LLMs has gained considerable attention within the machine learning community. This trend is evidenced by the Open LLM Leaderboard Beeching et al. (2023) data as of March 15th, 2024, which highlights the increasing prevalence of merged models among high-performing LLMs. Specifically, merged models represent 20% of the top 50 and 34% of the top 100 models, underscoring their significance in current LLM advancements. Plus, the current best-performing 3B555liminerity/Phigments12, 7B666liminerity/M7-7b, and the third top 13B [^7] RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2 open-source models on the Open LLM Leaderboard are merged models using MergeKit. In addition, the MergeKit repository’s meteoric rise in popularity is clearly illustrated by Figure 3, highlighting a significant and accelerating number of GitHub stars over time — an indicator of its growing influence and user endorsement in the developer community.
Figure 4 indicates a visual comparison of the performance of the top 50 open-source models ranked by their ‘Average’ scores on the Open LLM Leaderboard Park (2023). The ‘Average’ score is the mean value of evaluation scores from ‘ARC’Clark et al. (2018), ‘HellaSwag’Zellers et al. (2019), ‘MMLU’Hendrycks et al. (2021), ‘TruthfulQA’Lin et al. (2022), ‘Winogrande’Sakaguchi et al. (2019), ‘GSM8K’Cobbe et al. (2021) benchmarks. The dotted plot in black illustrates the parameter count for each model, and the bar chart represents the average score of each model. Merged models are depicted in ruby red bars, while the rest are colored in pink. The majority of leading merged models typically contain around 7 to 13 billion parameters. Among them, the top-performing merged model, ‘RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2’ [^8] https://huggingface.co/RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2, ranks 17th on the list. Remarkably, apart from models utilizing mixture of experts (MoE) Gale et al. (2023), majority of unmerged models surpassing this performance comprise approximately 35 to 70+ billion parameters. As depicted by the Open LLM Leaderboard, Merged models, particularly those incorporating mixture of experts (MoE) architecture, emerge as formidable contenders, offering superior performance while optimizing computational resources — a testament to the transformative potential of model merging in advancing the frontier of natural language processing.
@@ 6 Conclusion and Future Work
As stewards of the open-source community dedicated to the advancement of LLMs, our work with MergeKit underscores a commitment to democratizing access to cutting-edge AI technologies while fostering an environment of ethical integrity and continuous improvement. By providing an open-source toolkit that enables the merging of model checkpoints, we aim to enhance the collaborative capabilities of researchers, developers, and practitioners across the globe, encouraging innovation and the sharing of knowledge. In doing so, we are acutely aware of the necessity to uphold principles of fairness, accountability, and transparency within this community. This includes the proactive identification and mitigation of biases within merged models, ensuring the ethical use of data, and maintaining the privacy and security of information. Our commitment extends beyond technological advancements, encompassing the responsibility to engage with diverse stakeholders, gather feedback, and adapt our approaches to address ethical concerns effectively. We recognize the imperative to continually evolve our practices, striving for solutions that not only push the boundaries of AI but also do so with an unwavering commitment to the improvement of society.